An insurance company processes 800 claims per year. The average processing time for a claim is 5 weeks. 45 percent of all claims received are car insurance claims, 40 percent of all claims received are motorcycle insurance claims, 10 percent are boat insurance claims, and the remaining are house insurance claims. Hint: These are throughput values. On average there are, 20 car, 9 motorcycles, 12 boats, and some house claims in process. Hint: These are inventory values. Assume 50 weeks per year.

1. What is the average number of claims that are in process? The potential answers are:

A: 128 claims.

B: 64 claims.

C: 90 claims.

D: 80 claims.

E: 160 claims.

2. How many house insurance claims are in process? The potential answers are:

A: 77 claims.

B: 21 claims.

C: 72 claims.

D: 39 claims.

E: 45 claims.

3. How long, on average, does it take to process a car insurance claim? The potential answers are:

A: 7.5 weeks.

B: 4.5 weeks.

C: 3.75 weeks.

D: 6.67 weeks.

E: 2.78 weeks.

4. How long, on average, does it take to process a house insurance claim? The potential answers are:

A: 15.63 weeks.

B: 48.75 weeks.

C: 17.5 weeks.

D: 36 weeks.

E: 11.25 weeks.

please show work
Answer #1

1.According to little's law, WIP = lead time in weeks*units per week = 5*(800/50) = 80 claims

correct answer is D: 80 claims.

2.

Average number of house insurance claims in process = 80-20-9-12 = 39 claims

Correct answer is D. 39 claims

3.

Number of car insurance in process = 20

Units per week = (800/50)*45% = 7.2

So, according to little's law, time taken to process a car insurance claim = 20/7.2 = 2.777777778 = approximately 2.78


So correct answer is E. 2.78 weeks

4.
So, according to little's law, time taken to process a house insurance claim = WIP/weekly house insurance claims = 39/((800/50)*5%) = 48.75 weeks

Correct answer is B. 48.75 weeks